OpenStudy (celinegirl):

JUST CHECK MY ANSWER - A Determine the solution to f(x) = g(x) using the following system of equations: f(x) = 3x + 12 g(x) = −9.5x − 13 x = −2 x = −3 x = 4 x = 5

OpenStudy (er.mohd.amir):

3x+12=-9.5x-13 3x+9.5x=-13-12 12.5x=-25 x=-2

OpenStudy (celinegirl):

so A is correct?

OpenStudy (er.mohd.amir):

yes

OpenStudy (celinegirl):

Ok.. I solved it a different way but thank you!

OpenStudy (er.mohd.amir):

But ans is same if its correct.